To buy a computer system, a customer can choose one of 4 monitors, one of 2 keyboards, one of 4 computers and one of 3 printers. Determine the number of possible systems that a customer can choose from


1
Expert's answer
2022-03-31T07:23:09-0400

"4\cdot 2\cdot 4\cdot 3=96"